OpenStudy (anonymous):

consider factor group of Z4xZ6/<(0,1)> a. find elements in this group b. find all cosets Z4xZ6 with respect to <(0,1)>

OpenStudy (anonymous):

this should take a while \(\mathbb{Z}_4\times \mathbb{Z}_6\) has \(24\) elements so your factor group will have \(12\)

OpenStudy (anonymous):

well what are the cosets

OpenStudy (anonymous):

|Z4xZ6|=24 but we have to divide by element of <(0,1)> I am not sure how many elements does that have